A body starts from rest and accelerates uniformly at a rate of 5m/s2. calculate its velocity after 90m
1
Expert's answer
2019-03-11T12:17:20-0400
"v=v_0+at=at,\,\, s=\int v \, dt=\dfrac{at^2}{2}, \,\, t=\sqrt{\dfrac{2s}{a}}"

"v=\sqrt{2sa}=\sqrt{2\cdot 90 \cdot 5}=30 \, m/s"